# Interventional Time Series Priors for Causal Foundation Models

### 11 Mar 2026 v1

@@ 1 Introduction
We address this gap with CausalTimePrior [^1] https://github.com/thummd/CausalTimePrior, a framework for sampling temporal SCMs (TSCMs) together with paired observational and interventional time series (see Table 1). Our contributions are (1) a practical prior over discrete-time dynamic SCMs (Boeken and Mooij, 2024) that generates paired observational and interventional time series for training causal foundation models, (2) support for regime-switching TSCMs with Markov-driven structural breaks and interventional data—the first generator combining regime-switching dynamics (Balsells-Rodas et al., 2024) with interventional time series generation, and (3) preliminary experiments demonstrating that PFNs trained on CausalTimePrior learn causal structure from observational data alone.
@@ 3.3 Intervention Types
Given a sampled TSCM $\mathcal{S}$, we generate interventional data by modifying structural equations (Eberhardt and Scheines, 2007). Let $I\subseteq\{1,\ldots,N\}$ denote intervention targets and $t_{I}\subseteq\{1,\ldots,T\}$ the intervention times.
@@ 4 Experiments
We validate CausalTimePrior by analyzing a dataset of 100K generated TSCMs. (1) Structural diversity: the prior generates TSCMs with $N\in[3,10]$ variables, $K\in[1,3]$ lags, and $T=50$ time steps, including 70% diverse nonlinear TSCMs, 15% chain structures, and 15% regime-switching TSCMs. Erdős-Rényi sampling with varied edge probabilities implicitly produces canonical causal motifs (confounders, mediators, colliders) as subgraphs. (2) Stability: 0% divergence rate (no NaN/Inf values) across all 100K samples, achieved through value clipping and careful mechanism parameterization. (3) Intervention coverage: hard, soft, and time-varying interventions with mean effect size of 17.98 (std 53.93), demonstrating substantial variability in intervention magnitudes across types. (4) Paired data quality: observational and interventional series maintain similar statistics (obs: $\mu=46.78$, $\sigma=242.85$; int: $\mu=41.56$, $\sigma=228.52$), confirming that interventions produce realistic counterfactual outcomes rather than out-of-distribution artifacts. Example visualization and full distributions of prior properties are shown in Appendix B and C.
@@
As a preliminary demonstration, we train a simple 2-layer GRU-based PFN (128 hidden dim, 11 min on CPU) on 100K TSCMs from CausalTimePrior and evaluate on 1,000 held-out TSCMs. The model learns to distinguish causal from non-causal queries: Pred/GT ratio of 0.95 for intervened queries vs. 0.46 for non-causal queries (Table 2), and achieves comparable RMSE to per-dataset VAR baselines without per-sample fitting (Table 3). Implementation details, full results, baselines, a shuffled-intervention control experiment, ablations, generalizations, and and example are in Appendix D, E, F, G, and H.
@@ Appendix A CausalTimePrior Algorithm
Appendix B Example Visualizations
@@
Appendix C Prior Property Distributions
@@
Appendix D Implementation Details
@@ Appendix E Results
Query Type Queries RMSE $\downarrow$ NMSE $\downarrow$ Mean Pred Mean GT Pred/GT
@@
Appendix F Intervention Type Ablation
@@
Appendix G Out-of-Distribution Generalization
We evaluate whether the model generalizes to TSCMs with structural properties outside the training distribution. We generate an OOD test set of 1,000 TSCMs with: (1) larger graphs ($N\in[8,10]$ vs training mean $\sim$6), (2) maximum lag $K=3$, (3) denser graphs (edge probability $\geq 0.3$ vs training mean $\sim$0.29), and (4) only complex nonlinear mechanisms (sin, cos, square, tanhReLU).
@@ Appendix G Out-of-Distribution Generalization
Appendix H Causal vs. Correlation-Based Prediction

### 08 Apr 2026 v2

@@ 1 Introduction
We address this gap with CausalTimePrior [^1] https://github.com/thummd/CausalTimePrior, a framework for sampling temporal SCMs (TSCMs) together with paired observational and interventional time series (see Table 1). Our contributions are (1) a practical prior over discrete-time dynamic SCMs (Boeken and Mooij, 2024) that generates paired observational and interventional time series for training causal foundation models, (2) support for regime-switching TSCMs with Markov-driven structural breaks and interventional data—the first generator combining regime-switching dynamics (Balsells-Rodas et al., 2024) with interventional time series generation, and (3) preliminary experiments demonstrating that PFNs trained on CausalTimePrior can predict interventional outcomes given only observational context.
@@ 3.3 Intervention Types
Given a sampled TSCM $\mathcal{S}$, we generate interventional data by modifying structural equations (Eberhardt and Scheines, 2007). Let $I\subseteq\{1,\ldots,N\}$ denote intervention targets and $t_{I}\subseteq\{1,\ldots,T\}$ the intervention times. We provide examples for each intervention type in Appendix C.
@@ 3.5 Regime-Switching Priors
3.6 Design Choices and Assumptions
Complexity, identifiability, and causality.
[…]
Noise assumptions and prediction horizon.
Noise is additive and exogenous: $\epsilon_{t}^{(i)}$ is added after the nonlinear activation and sampled independently of parent values. It is also Markovian—i.i.d. across time—implying that the causal graph fully mediates all temporal dependencies; extending to non-Markovian (temporally correlated) noise would model latent confounders persisting across time and is an important future direction. At training time, the PFN queries outcomes 0–5 steps after intervention onset (70% downstream queries at 1–5 steps, 30% instantaneous), concentrating learning on the short-range causal propagation window; see Appendix B for further discussion.
@@ 4 Experiments
We validate CausalTimePrior by analyzing a dataset of 100K generated TSCMs. (1) Structural diversity: the prior generates TSCMs with $N\in[3,10]$ variables, $K\in[1,3]$ lags, and $T=50$ time steps, including 70% diverse nonlinear TSCMs, 15% chain structures, and 15% regime-switching TSCMs. Erdős-Rényi sampling with varied edge probabilities implicitly produces canonical causal motifs (confounders, mediators, colliders) as subgraphs. (2) Stability: 0% divergence rate (no NaN/Inf values) across all 100K samples, achieved through value clipping and careful mechanism parameterization. (3) Intervention coverage: hard, soft, and time-varying interventions with mean effect size of 17.98 (std 53.93), demonstrating substantial variability in intervention magnitudes across types. (4) Paired data quality: observational and interventional series maintain similar statistics (obs: $\mu=46.78$, $\sigma=242.85$; int: $\mu=41.56$, $\sigma=228.52$), confirming that interventions produce realistic counterfactual outcomes rather than out-of-distribution artifacts. Example visualization and full distributions of prior properties are shown in Appendix D and E.
@@
As a preliminary demonstration, we train a simple 2-layer GRU-based PFN (128 hidden dim, 11 min on CPU) on 100K TSCMs from CausalTimePrior and evaluate on 1,000 held-out TSCMs. The model learns to distinguish causal from non-causal queries: Pred/GT ratio of 0.95 for intervened queries vs. 0.46 for non-causal queries (Table 2), and achieves comparable RMSE to per-dataset VAR baselines without per-sample fitting (Table 3). Implementation details, full results, baselines, a shuffled-intervention control experiment, ablations, generalizations, and an example are in Appendix F, G, H, I, and J.
@@ Appendix A CausalTimePrior Algorithm
Appendix B Prior Assumptions and Limitations
This section details the modeling assumptions underlying CausalTimePrior and their implications for identifiability and generalization.
Effect propagation and decay.
Causal effects in the prior decay through three mechanisms: (1) the maximum lag $K\in\{1,2,3\}$ bounds direct causal influence to at most 3 steps back, (2) geometric edge decay ($p_{k}=p\cdot 0.7^{k}$) makes distant lag connections increasingly sparse, and (3) bounded activations (e.g., $\tanh$, $\sin$) naturally dampen signal propagation. There is no explicit spectral radius constraint on the weight matrices; instead, stability is enforced empirically via value clipping to [1000,1000][-1000,1000] at each simulation step and divergence rejection (any trajectory with $|x|>10^{6}$ or NaN/Inf is discarded and resampled). A 50-step burn-in period is discarded to approximate the stationary distribution.
This design implies that causal effects attenuate rapidly—typically within 1–5 steps after intervention onset. While this is appropriate for systems with short-range temporal dependencies, it may underweight long-range causal effects (e.g., economic policy impacts over months). A PFN trained on this prior would inherit this short-range bias.
Markovian noise and identifiability.
The noise $\epsilon_{t}^{(i)}$ is sampled i.i.d. at each time step—independent across both time and variables. This Markovian assumption implies that the causal graph $\mathcal{G}$ fully mediates all temporal dependencies: there are no hidden common causes acting across time beyond what the lagged edges encode. Identifiability within the modeled lag window follows from the additive noise model (ANM) structure (Peters et al., 2014): each mechanism has the form $\sigma(\sum w_{j}x_{j}+b)+\epsilon$, and the use of non-Gaussian noise families (Uniform, Laplace) for non-root nodes supports distinguishing causal from non-causal associations under standard ANM identifiability results.
An extension to non-Markovian confounding—where $\text{Cov}(\epsilon_{t}^{(i)},\epsilon_{t-s}^{(j)})\neq 0$ for some s>0s>0—would model latent confounders that persist over time (e.g., unobserved trends or regime states not captured by the graph). This would break the assumption that interventional distributions depend only on the explicit causal structure and is an important direction for increasing the prior’s realism.
Interventional vs. counterfactual semantics.
CausalTimePrior generates interventional paired data, not counterfactual paired data. Concretely, when generate_pair produces an observational series $\mathbf{X}^{\text{obs}}$ and an interventional series $\mathbf{X}^{\text{int}}$, the two simulations draw independent noise realizations $\{\epsilon_{t}^{(i)}\}$ and $\{\tilde{\epsilon}_{t}^{(i)}\}$. This means the trajectories may differ even outside the intervention window—not because of any causal effect, but because they are driven by different exogenous noise.
This corresponds to the interventional query $P(\mathbf{X}\mid\text{do}(X^{(i)}_{t_{I}}=c))$: “what would a new draw from the system look like under this intervention?” The alternative is the counterfactual query $P(\mathbf{X}^{CF}\mid\mathbf{X}^{\text{obs}},\text{do}(X^{(i)}_{t_{I}}=c))$: “given this specific observational realization, what would have happened if we had intervened?” Counterfactual inference requires the three-step abduction-action-prediction procedure (Pearl, 2009): (1) infer the noise $\epsilon=\mathbf{X}^{\text{obs}}-f(\text{Pa})$ (abduction), (2) modify the structural equation for the intervened variable (action), and (3) re-simulate with the same noise (prediction).
In our framework, generating counterfactual pairs would require sharing the pre-sampled noise tensor between the observational and interventional simulation runs, so that the two trajectories are identical before the intervention onset and diverge only through causal propagation of the intervention effect. This is a natural extension that would enable training PFNs for counterfactual estimation—a strictly harder task than interventional prediction, since it requires reasoning about unit-level rather than population-level effects.
Noise model details.
All noise is additive (added after the nonlinear activation) and exogenous (independent of parent values). Root nodes (no parents in $G_{0}$) receive Gaussian noise with $\text{std}\sim\text{ShiftedExp}(\text{rate}=1.0,\text{shift}=0.1)$, providing larger driving noise. Non-root nodes receive noise from one of three variance-matched families (each with probability $\tfrac{1}{3}$): Gaussian $\mathcal{N}(0,\text{std}^{2})$, Uniform U(a,a)U(-a,a) with $a=\text{std}\sqrt{3}$, or Laplace $\text{Lap}(0,b)$ with $b=\text{std}/\sqrt{2}$, where $\text{std}\sim\text{ShiftedExp}(\text{rate}=10.0,\text{shift}=0.01)$. The smaller non-root noise ensures that most of a variable’s variance comes from its causal parents rather than its own noise term.
Stationarity scope.
Within a single SCM instance, mechanisms are stationary: weights, biases, activation functions, and noise distributions are fixed across all time steps. Non-stationarity enters only through regime-switching SCMs (15% of samples), where 2–3 regimes—each with its own causal graph and mechanisms—alternate according to a sticky Markov chain. The current framework does not support continuously time-varying coefficients, one-time structural breaks, or heteroscedastic noise. These extensions would broaden the prior’s coverage of real-world non-stationary dynamics.
PFN prediction horizon.
During training, 70% of queries target a different variable 1–5 steps after intervention onset (downstream causal effects), while 30% query the intervention target itself at the intervention time (instantaneous effects). The PFN architecture accepts arbitrary $(\text{intervention\_time},\text{query\_time})$ pairs as normalized inputs and can in principle predict at any horizon, but accuracy is expected to degrade beyond the 0–5 step training concentration. Extending the training distribution to longer horizons—potentially with curriculum learning (Bengio et al., 2009)—could improve long-range causal effect estimation.
Appendix C Intervention Type Examples
Figure 1 illustrates the three intervention types supported by CausalTimePrior, applied to the same 10-variable TSCM. Each row shows one type: (1) Hard intervention ($\text{do}(X_{0}:=3.0)$) replaces the variable’s mechanism entirely with a constant, severing all incoming causal edges during the intervention window—visible as a flat interventional trajectory. (2) Soft intervention ($\delta=1.5$) adds an additive shift to the mechanism output while preserving parental influence—the interventional trajectory tracks the observational one but is displaced. (3) Time-varying intervention ($\text{do}(X_{0}:=2\sin(2\pi t/L))$) replaces the mechanism with a time-dependent function, producing a sinusoidal interventional trajectory.
Figure 1: The three intervention types in CausalTimePrior. Left column: the intervention signal applied to $X_{0}$. Right column: resulting observational (blue) vs. interventional (red) trajectories. Yellow shading marks the intervention window $[30,70)$.
Figure 2 shows the four time-varying intervention profiles in detail, each applied to the same SCM and variable. The step profile jumps from 2-2 to +2+2 at the midpoint; the ramp linearly interpolates from 2-2 to +2+2; the sinusoidal profile follows $c(t)=2\sin(2\pi t/L)$; and the sampled trajectory draws independent $c(t)\sim\mathcal{N}(0,4)$ at each time step. The dashed line shows the intervention signal $c(t)$ on the secondary axis.
Figure 2: Four time-varying intervention profiles. Blue: observational trajectory. Red: interventional trajectory. Dashed: intervention signal $c(t)$ (right axis). All four profiles are applied to the same variable in the same TSCM.
Appendix D Example Visualizations
@@
Appendix E Prior Property Distributions
@@
Appendix F Implementation Details
@@ Appendix G Results
Query Type Queries RMSE $\downarrow$ NMSE $\downarrow$ Mean Pred Mean GT Pred/GT $\uparrow\downarrow=1$
@@
Appendix H Intervention Type Ablation
@@
Appendix I Out-of-Distribution Generalization
We evaluate whether the model generalizes to TSCMs with structural properties outside the training distribution. We generate an out-of-distribution (OOD) test set of 1,000 TSCMs with: (1) larger graphs ($N\in[8,10]$ vs training mean $\sim$6), (2) maximum lag $K=3$, (3) denser graphs (edge probability $\geq 0.3$ vs training mean $\sim$0.29), and (4) only complex nonlinear mechanisms (sin, cos, square, tanhReLU).
@@ Appendix I Out-of-Distribution Generalization
Appendix J Causal vs. Correlation-Based Prediction